# PPFPL: Cross-silo Privacy-preserving Federated Prototype Learning Against Data Poisoning Attacks on Non-IID Data

### v4

@@ IV Problem Statement
IV-A PPFPL framework
The framework of PPFPL consists of four entities, each of which has its specific function, as shown in Fig. 1. The interaction among entities composes operation of whole system. The specific functions of each entity are outlined as follows.
Figure 1: The PPFPL framework. ① KGC generates $Pkv/Skv$ for Verifier and $Pkx/Skx$ for Clients. ② After local training is completed, clients submit encrypted local prototypes to the Aggregator. ③ Verifier and Aggregator perform secure aggregation protocol to get encrypted global prototypes. ④ Aggregator distributes encrypted global prototypes to Clients.
  • Key Generation Center (KGC). The entity is responsible for generating and managing keys of both Clients and Verifier, which are important elements to ensure security of encryption/decryption process.
  • Clients. The Clients are large organizations participating in federated training. The aim of benign organizations is to get a better model by federated training. They has a pair of public/secret keys generated by KGC, denoted as $Pkx$/$Skx$.
@@ IV-A PPFPL framework
  • Verifier. The Verifier is a non-colluding central server and calculates cooperatively with Aggregator to aggregate local prototypes. It has a pair of public/private keys generated by KGC, denoted as $Pkv$/$Skv$.
@@
IV-B Potential Threats for PPFPL
@@
ii) The Aggregator and Verifier are considered as “non-colluding” and “curious but honest”. Specifically, we assume that Aggregator and Verifier do not collude to attack PPFPL. The assumption is reasonable in practice, since it is usually impossible for two well-known service providers to collude with each other due to legal regulations, company reputation [47]. Furthermore, we assume that they follow the system’s protocol but may attempt to get sensitive information (i.e., raw training data) by inferring client-submitted local prototypes. Notably, although prototype is different from gradient, traditional inference attack methods cannot recover client’s private training data from prototype. However, the work in [48] designs a dynamic memory model inversion attack that can recover the private training data by utilizing client’s learned prototypes. Thus, preserving the privacy of prototypes is necessary. In addition, we do not consider security attacks resulting from employee insider threats (e.g., compromised employees within the Aggregator or Verifier).
iii) In practice, we cannot guarantee that all clients in the system are honest. Therefore, clients can be either benign or malicious. Specifically, benign clients are “honest but curious”. For malicious clients, they can collude together to infer sensitive information about benign clients. We consider that the proportion of malicious clients is less than 50%, which is a more realistic threat in cross-silo FL due to the high reputation of large organizations participating. Furthermore, malicious clients can launch data poisoning attacks, and submit malicious local prototypes derived from the poisoned data. In federated prototype learning, we define the following two representative types of data poisoning attacks from the perspective of data features and labels.
@@
IV-C Design Goals of PPFPL
The goal of our study is to improve performance in cross-silo PPFL under poisoned Non-IID data while resisting data poisoning attacks. Specifically, we design the PPFPL framework to fulfill the following goals:
  • Security. PPFPL should guarantee the correct model training in the presence of data poisoning attacks with Non-ID data. In other words, the model performance of each benign client is not affected by data poisoning attacks under different data distributions.
  • Privacy. PPFPL should ensure privacy and security of benign clients. For any third entity, they cannot access the sensitive information about benign clients.
  • Efficiency. PPFPL should reduce the number of parameters submitted to two servers compared to other similar frameworks, thereby reducing privacy computation overheads and communication overheads.
V Design of PPFPL
@@ V-A Overview of PPFPL
  • Step I. Local Computation: Each client trains its local model with its local dataset. Then, the client normalizes and encrypts its local prototypes and submits them to Aggregator.
  • Step II. Secure Aggregation Protocol: The two servers verify the normalization of encrypted local prototypes submitted by clients, and performs the secure two-party computation to get encrypted global prototypes, and distribute them to each client.
The above process is stopped until configured number of communication rounds $T$. In the following, we describe the process of Steps I and II in detail.
@@ 13       Aggregator distributes global prototypes to each client;
V-B Local Computation
The local computation step includes two essential stages: local model training and prototype handling stages. The detail is outlined in Algorithm 2.
@@ V-B Local Computation
1 for each client $m$in $\mathcal{S}$ do
@@
3       Decrypt $\{\llbracket\textit{{C}}^{k}_{t}\rrbracket_{Pkx}\}_{k\in\mathcal{I}}^{|\mathcal{I}|}$ using its $Skx$;
@@
6       Send $\{\llbracket\widetilde{\textit{{c}}}_{m,t}^{k}\rrbracket_{Pkv}\}_{k\in\mathcal{I}}^{|\mathcal{I}|}$ to Aggregator;
@@
V-B1 Local Model Training
During local model training, each client aims to minimize classification loss while aligning its local prototype close to the global prototype. To achieve this, we design an auxiliary term by leveraging prototype learning in the local optimization function. Formally, the local optimization function of client $m$ is defined as:
$\mathcal{L}(\textit{{w}}_{m};\mathcal{D}$ ,${}_{m},\textit{{c}}^{k}_{m},\textit{{C}}^{k})=\mathcal{L_{S}}(\mathcal{F}_{m}(\textit{{w}}_{m};\textit{{x}}_{(i)}),y_{(i)})$ (3)
$+\lambda\mathcal{L_{R}}(\textit{{c}}^{k}_{m},\textit{{C}}^{k}),\forall(\textit{{x}}_{(i)},y_{(i)})\in\mathcal{D}_{m},\forall k\in\mathcal{I},$
where $\mathcal{L_{S}}(\cdot,\cdot)$ is the classification loss function (e.g., cross-entropy loss function), $\lambda$ is the importance weight of auxiliary term, and $\mathcal{L_{R}}(\cdot,\cdot)$ is the auxiliary term, defined as
$\mathcal{L_{R}}(\textit{{c}}^{k}_{m},\textit{{C}}^{k})=\frac{1}{|\mathcal{I}|}\sum_{k\in\mathcal{I}}(1-\textit{sim}(\textit{{c}}_{m}^{k},\textit{{C}}^{k})),$ (4)
where $\textit{sim}(\cdot,\cdot)$ denotes cosine similarity between two vectors. The cosine similarity ranges from -1 to 1, where the value closer to “1” indicates similar vector directions, and conversely, the value closer to “-1” means opposing directions. The optimization function ensures that each client reduces the classification loss while aligning its local prototype with the global prototype in direction. Next, we present the process of local model training.
Specifically, in the $t$-th communication round, each client $m$ uses its local dataset $\mathcal{D}_{m}$ to iteratively train its local model. This iterative process is executed in Algorithm 3. The superscript $(e)$ indicates the iteration state of variables, where $e\in\{1,\cdots,E\}$. Each client $m$ uses its local model parameters $\textit{{w}}_{m,t-1}^{(E)}$ from the $(t-1)$-th round as the starting point in the current $t$-th round. Subsequently, each client iteratively performs the following stages.
  • In $e$-th local iteration, the client randomly selects training data $\mathcal{D}_{m}^{(e)}$ from its local dataset $\mathcal{D}_{m}$.
@@ V-B1 Local Model Training
    where $\nabla$ denotes derivation operation. Then, the local model $\textit{{w}}_{m,t}^{(e-1)}$ is updated via the following formula:
@@
After $E$ local iterations of the above process, each client derives the updated local model parameters $\textit{{w}}_{m,t}^{(E)}$.
@@ 1 $\textit{{w}}_{m,t}^{(0)}=\textit{{w}}_{m,t-1}^{(E)}$;
2 for each local iteration $e$in $\{1,2,\cdots,E\}$ do
@@ 3       Randomly sample $\mathcal{D}_{m}^{(e)}\subset\mathcal{D}_{m}$;
4       for each class $k$in $\mathcal{I}$ do
@@ V-B2 Prototype Handling
Since the parameters of local model change with each iteration, the prototypes generated by each local iteration are different. Consequently, the local prototypes evolve dynamically during local model training. To submit more representative prototypes, each client regenerates them via the local model parameters $\textit{{w}}_{m,t}^{(E)}$. Formally, let $\textit{{c}}^{k}_{m,t}$ be the submitted local prototype, is computed as:
$\textit{{c}}^{k}_{m,t}=\frac{1}{|\mathcal{D}^{k}_{m}|}\sum_{(\textit{{x}}_{(i)},y_{(i)})\in\mathcal{D}^{k}_{m}}f_{m}(\textit{{r}}_{m,t}^{(E)};\textit{{x}}_{(i)}),\forall k\in\mathcal{I},$
@@
Considering that malicious clients amplify their submitted local prototypes, the local prototype of each class in each client is normalized. Formally, the local prototype is normalized by the following formula:
@@
where $\widetilde{\textit{{c}}}^{k}_{m,t}$ is a unit vector. After normalization, each local prototype needs to be encrypted for privacy protection.
@@
1 for each class $k$in $\mathcal{I}$ do
@@
4       Encrypt $\widetilde{\textit{{c}}}^{k}_{m,t}$ by Verifier’s $Pkv$ to get $\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}$;
@@
To protect privacy of clients, they encrypts their own normalized local prototypes $\widetilde{\textit{{c}}}^{k}_{m,t}$ using the Verifier’s public key $Pkv$ to get $\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}$, and sends the $\{\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}\}_{k\in\mathcal{I}}^{|\mathcal{I}|}$ to Aggregator, where CKKS technique is used to encrypt.
V-C Secure Aggregation
@@
V-C1 Normalization verification
Since malicious clients may bypass the normalization stage to amplify the impact of their local prototypes, the two servers need to verify that encrypted local prototypes are normalized. Specifically, Aggregator calculates the inner product $\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}\cdot\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}$ for each local prototype, and sends the results to Verifier, where $\cdot$ denotes the inner product. Then, the Verifier decrypts $\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}\cdot\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}$ using its secret key $Skv$, and checks whether the inner product $\|\widetilde{\textit{{c}}}^{k}_{m,t}\|^{2}$ equals 1. If the inner product of prototype from client $m$ is not equal to 1, which indicates that its local prototype does not normalized, client $m$ is removed from the set of clients $\mathcal{S}$. After validation, Verifier sends the remaining client set $\mathcal{S}$ to Aggregator.
V-C2 Secure Two-party Computation
To achieve privacy preservation and Byzantine-robust aggregation results, two servers perform secure two-party computation to calculate global prototypes. The computing of global prototype is shown in Algorithm 5 and described as follows.
@@ 6       Randomly select a $n$-dimensional vector $\textit{{V}}^{n}$ and a number $p$;
10      Decrypt $p\times\llbracket h^{k}_{m,t}\rrbracket_{Pkv}$and $\textit{{V}}^{n}\odot\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}$ with $Skv$;
@@ 12       Compute $\textit{Sum}^{k}_{t}$;
[…]
@@ 19       Distribute global prototype $\llbracket\textit{{C}}^{k}_{t+1}\rrbracket_{Pkx}$ to clients;
$\llbracket\textit{{C}}^{\prime k}_{t+1}\rrbracket_{Pkv}$ $=\frac{1}{|\mathcal{S}|}\sum_{m\in\mathcal{S}}\llbracket\textit{{c}}_{m,t}^{k}\rrbracket_{Pkv},\forall k\in\mathcal{I},$ (5)
where $\llbracket\textit{{C}}^{\prime k}_{t+1}\rrbracket_{Pkv}$ denotes the trusted prototype for class $k$ in $(t+1)$-th communication round. $\textit{{C}}^{\prime k}_{t+1}$ is considered a trusted prototype for two reasons: 1) In terms of magnitude, this is because the normalized malicious prototype do not affect the magnitude of trusted prototype, only the direction of trusted prototype; 2) In terms of direction, the average of all prototypes remains a plausible direction due to the small proportion of malicious clients. Subsequently, Aggregator obtains the plaintext $\|\textit{{C}}^{\prime k}_{t+1}\|$ for subsequent cosine similarity computation. To achieve the plaintext, it computes the inner product $\llbracket\textit{{C}}^{\prime k}_{t+1}\rrbracket_{Pkv}\cdot\llbracket\textit{{C}}^{\prime k}_{t+1}\rrbracket_{Pkv}$ and sends it to the Verifier for decryption, then receives the decrypted result. Then, cosine similarity, a widely used measure of the angle between two vectors, is employed to measure the credibility of local prototype. If the direction of local prototype is similar to that of trusted prototype, its credibility is higher. Formally, Aggregator computes the credibility of local prototype by the following formula,
$\llbracket\textit{sim}^{k}_{m,t}\rrbracket_{Pkv}$ $=\llbracket\textit{sim}(\widetilde{\textit{{c}}}^{k}_{m,t},\textit{{C}}^{\prime k}_{t+1})\rrbracket_{Pkv}=\left\llbracket\frac{\widetilde{\textit{{c}}}^{k}_{m,t}\cdot\textit{{C}}^{\prime k}_{t+1}}{\|\widetilde{\textit{{c}}}^{k}_{m,t}\|\|\textit{{C}}^{\prime k}_{t+1}\|}\right\rrbracket_{Pkv}$ (6)
$=\frac{1}{\|\textit{{C}}^{\prime k}_{t+1}\|}\left(\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}\cdot\llbracket\textit{{C}}^{\prime k}_{t+1}\rrbracket_{Pkv}\right),$
where $\llbracket\textit{sim}^{k}_{m,t}\rrbracket_{Pkv}$ denotes the credibility of local prototype. Additionally, we set the detection threshold $\chi$ and define local prototypes with its credibility less than $\chi$ as anomalous local prototypes. During global prototype aggregation, we set the aggregation weight of anomalous local prototype to 0. Formally, the aggregation weight of each local prototype is calculated as follows:
@@
where $j^{k}_{m,t}$ denotes the aggregation weight of local prototype for class $k$ at client $m$ in the $t$-th communication round. However, Aggregator cannot directly compare $\llbracket\textit{sim}^{k}_{m,t}\rrbracket_{Pkv}$ with $\chi$ under encryption domain.
To implement comparison within encryption domain, the work in [49] offers a ciphertext comparison method that outputs the maximum value of homomorphic ciphertexts corresponding to two plaintexts in the range (0, 1) without decryption. However, since the cosine similarity ranges from -1 to 1, and the method cannot be applied directly. To resolve the range mismatch, Aggregator sets $\llbracket\textit{sim}^{\prime k}_{m,t}\rrbracket=\frac{1}{2}(\llbracket\textit{sim}^{k}_{m,t}\rrbracket+\llbracket 1\rrbracket)$ and $\chi^{\prime}=\frac{1}{2}(\chi+1)$ as input to the comparison method [49]. This is because the size relationship between $\llbracket\textit{sim}^{k}_{m,t}\rrbracket$ and $\chi$ remains unchanged. Then, Aggregator encrypts $\chi^{\prime}$ to get $\llbracket\chi^{\prime}\rrbracket$, and calculates the maximum value $\llbracket h^{k}_{m,t}\rrbracket$ between $\llbracket\textit{sim}^{\prime k}_{m,t}\rrbracket$ and $\llbracket\chi^{\prime}\rrbracket$ by Algorithm 6.
@@
1 $a=\llbracket\textit{sim}^{\prime k}_{m,t}\rrbracket,b=\llbracket\chi^{\prime}\rrbracket$;

### v5

@@ 4 Problem Statement
4.1 PPFPL framework
The framework of PPFPL consists of four entities, each of which has its specific function, as shown in Fig. 1. The interaction among these entities composes the operation of whole system. The roles of each entity are outlined as follows.
Figure 1: The PPFPL framework. ① KGC generates $Pkv/Skv$ for Verifier and $Pkx/Skx$ for Clients. ② After local training is completed, clients submit encrypted local prototypes to the Aggregator. ③ Verifier and Aggregator perform a secure aggregation protocol to get encrypted global prototypes. ④ Aggregator distributes the encrypted global prototypes to Clients.
  • Key Generation Center (KGC). The entity is responsible for generating and managing the keys of both Clients and Verifier, which are essential elements to ensure security of encryption/decryption process.
  • Clients. The Clients are large organizations participating in federated training. The aim of benign organizations is to get a better model by federated training. They have a pair of public/secret keys generated by KGC, denoted as PkxPkx/SkxSkx.
@@ 4.1 PPFPL framework
  • Verifier. The Verifier is a non-colluding central server that cooperates with the Aggregator to aggregate local prototypes. It has a pair of public/private keys generated by KGC, denoted as PkvPkv/SkvSkv.
@@
4.2 Potential Threats for PPFPL
@@
ii) The Aggregator and Verifier are assumed to be non-colluding and curious but honest. Specifically, we assume that Aggregator and Verifier do not collude to attack PPFPL. The assumption is reasonable in practice, since it is usually impossible for two well-known service providers to collude with each other due to legal regulations and company reputation [50]. Furthermore, we assume that they follow the system’s protocol but may attempt to get sensitive information (i.e., raw training data) by inferring client-submitted local prototypes. Notably, although prototypes differ from gradients, traditional inference attack methods cannot recover client’s private training data from prototype. Although prototypes differ from gradients, and traditional inference attacks cannot directly reconstruct private training data from them, the work in [44] designs a dynamic memory model inversion attack that can recover the private training data by utilizing client’s learned prototypes. Thus, preserving the privacy of prototypes is necessary. Notably, we do not consider security attacks resulting from employee insider threats (e.g., compromised employees within the Aggregator or Verifier).
iii) In practice, we cannot guarantee that all clients in the system are honest. Therefore, clients can be either benign or malicious. Specifically, benign clients are “honest but curious”. For malicious clients, they can collude to infer sensitive information about benign clients. We assume that the proportion of malicious clients is less than 50%, which is a more realistic threat in cross-silo FL due to the high reputation of large organizations participating. Furthermore, malicious clients can launch data poisoning attacks by submitting malicious prototypes derived from poisoned data. In federated prototype learning, we define the following two representative types of data poisoning attacks from the perspective of data features and labels.
@@
4.3 Design Goals of PPFPL
The goal of our study is to improve model performance in cross-silo federated learning under poisoned Non-IID data while resisting data poisoning attacks. Specifically, we design the PPFPL framework to fulfill the following goals:
  • Security. PPFPL should guarantee the robust model training in the presence of data poisoning attacks with Non-IID data. In other words, the model performance of each benign client is not affected by data poisoning attacks under different data distributions.
  • Privacy. PPFPL should ensure the privacy and security of benign clients. For any third entity, they cannot access the sensitive information about benign clients.
  • Efficiency. PPFPL should reduce the number of parameters submitted to two servers compared to other similar schemes, thereby reducing both privacy computation and communication overheads.
5 Design of PPFPL
@@ 5.1 Overview of PPFPL
  • Step I. Local Computation: Each client trains its local model with its local dataset. Then, the client normalizes and encrypts its local prototypes before submitting them to Aggregator.
  • Step II. Secure Aggregation Protocol: The two servers verify the normalization of encrypted local prototypes submitted by clients, and perform the secure two-party computation to get encrypted global prototypes, which are then distributed to each client.
The above process repeats until configured number of communication rounds $T$. In the following, we describe the process of Steps I and II in detail.
@@ 13  Aggregator distributes global prototypes to each client;
5.2 Local Computation
The local computation step includes two essential stages: local model training and prototype handling stages. The details are outlined in Algorithm 2.
@@ 5.2 Local Computation
1 for each client $m$ $\in$ $\mathcal{S}$ do
@@
3  Decrypt $\{\llbracket\textit{{C}}^{k}_{t}\rrbracket_{Pkx}\}_{k\in\mathcal{I}}^{|\mathcal{I}|}$ using its SkxSkx;
@@
6  Send
$\{\llbracket\widetilde{\textit{{c}}}_{m,t}^{k}\rrbracket_{Pkv}\}_{k\in\mathcal{I}}^{|\mathcal{I}|}$
to Aggregator;
@@ 6  Send
5.2.1 Local Model Training
During local model training, each client aims to minimize its classification loss while aligning its local prototype close to the global prototype. To achieve this, we design an auxiliary term by leveraging prototype learning in the local optimization function. Formally, the local optimization function of client $m$ is defined as:
$\begin{aligned} \mathcal{L}(\textit{{w}}_{m};\mathcal{D}&{}_{m},\textit{{c}}^{k}_{m},\textit{{C}}^{k})=\mathcal{L_{S}}(\mathcal{F}_{m}(\textit{{w}}_{m};\textit{{x}}_{(i)}),y_{(i)})\\ &+\lambda\mathcal{L_{R}}(\textit{{c}}^{k}_{m},\textit{{C}}^{k}),\forall(\textit{{x}}_{(i)},y_{(i)})\in\mathcal{D}_{m},\forall k\in\mathcal{I},\end{aligned}$
where $\mathcal{L_{S}}(\cdot,\cdot)$ is the classification loss function (e.g., cross-entropy loss function), $\lambda$ is the importance weight of the auxiliary term, and $\mathcal{L_{R}}(\cdot,\cdot)$ is the auxiliary term, defined as:
$\mathcal{L_{R}}(\textit{{c}}^{k}_{m},\textit{{C}}^{k})=\frac{1}{|\mathcal{I}|}\sum_{k\in\mathcal{I}}(1-\textit{sim}(\textit{{c}}_{m}^{k},\textit{{C}}^{k})),$
where $\textit{sim}(\cdot,\cdot)$ denotes cosine similarity between two vectors. The cosine similarity ranges from 1-1 to $1$, where the value closer to 1``1" indicates similar vector directions, while value closer to 1``-1" means opposing directions. The optimization function ensures that each client reduces its classification loss while aligning its local prototype with the global prototype in direction. Next, we present the process of local model training.
Specifically, in the $t$-th communication round, each client $m$ uses its local dataset $\mathcal{D}_{m}$ to iteratively train its local model, as detailed in Algorithm 3. The superscript $(e)$ indicates the iteration state of variables, where $e\in\{1,\cdots,E\}$. Each client $m$ uses its local model parameters $\textit{{w}}_{m,t-1}^{(E)}$ from the (t1)(t-1)-th round as the starting point in the current $t$-th round. Subsequently, each client iteratively performs the following stages.
  • In the $e$-th local iteration, the client randomly selects training data $\mathcal{D}_{m}^{(e)}$ from its local dataset $\mathcal{D}_{m}$.
@@ 5.2.1 Local Model Training
    where $\nabla$ denotes derivation operation. The local model parameters $\textit{{w}}_{m,t}^{(e-1)}$ are then updated as:
@@
After completing $E$ local iterations, each client derives the updated local model parameters $\textit{{w}}_{m,t}^{(E)}$.
@@ 1 $\textit{{w}}_{m,t}^{(0)}=\textit{{w}}_{m,t-1}^{(E)}$;
2 for each local iteration $e$ $\in$ $\{1,2,\cdots,E\}$ do
@@ 3  Randomly sample $\mathcal{D}_{m}^{(e)}\subset\mathcal{D}_{m}$;
4  for each class $k$ $\in$ $\mathcal{I}$ do
@@ 5.2.2 Prototype Handling
Since the parameters of local model change with each iteration, the prototypes generated by each local iteration are different. Consequently, the local prototypes evolve dynamically during local model training. To submit more representative prototypes, each client regenerates them via the local model parameters $\textit{{w}}_{m,t}^{(E)}$ . Formally, let $\textit{{c}}^{k}_{m,t}$ be the submitted local prototype, which is computed as:
$\begin{aligned} \textit{{c}}^{k}_{m,t}=\frac{1}{|\mathcal{D}^{k}_{m}|}\sum_{(\textit{{x}}_{(i)},y_{(i)})\in\mathcal{D}^{k}_{m}}f_{m}(\textit{{r}}_{m,t}^{(E)};\textit{{x}}_{(i)}),\forall k\in\mathcal{I},\end{aligned}$
@@
Considering that malicious clients amplify their submitted local prototypes, the local prototype of each class at each client is normalized. Formally, the normalized local prototype is given by:
@@
where $\widetilde{\textit{{c}}}^{k}_{m,t}$ is a unit vector. After normalization, each local prototype is encrypted to ensure privacy protection.
@@
1 for each class $k$ $\in$ $\mathcal{I}$ do
@@
4  Encrypt $\widetilde{\textit{{c}}}^{k}_{m,t}$ by Verifier’s PkvPkv to get $\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}$;
@@
To protect privacy of clients, they encrypt their own normalized local prototypes $\widetilde{\textit{{c}}}^{k}_{m,t}$ using the Verifier’s public key $Pkv$ to get $\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}$ , and sends the $\{\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}\}_{k\in\mathcal{I}}^{|\mathcal{I}|}$ to Aggregator, where CKKS technique is used to encrypt.
5.3 Secure Aggregation Protocol
@@
5.3.1 Normalization Verification
Since malicious clients may amplify the impact of their local prototypes, the two servers need to verify that encrypted local prototypes are normalized. Specifically, Aggregator calculates the inner product $\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}\cdot\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}$ for each local prototype, and sends the results to Verifier, where $\cdot$ denotes the inner product. Then, the Verifier decrypts $\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}\cdot\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}$ using its secret key $Skv$, and checks whether the inner product $\|\widetilde{\textit{{c}}}^{k}_{m,t}\|^{2}$ equals 1. If the inner product of prototype from client $m$ is not equal to 1, which indicates that its local prototype is not normalized, client $m$ is removed from the set of clients $\mathcal{S}$. After validation, Verifier sends the remaining client set $\mathcal{S}$ to Aggregator.
5.3.2 Secure Two-party Computation
To ensure privacy preservation and Byzantine-robust aggregation, two servers perform secure two-party computation to calculate global prototypes. The computation of the global prototypes is shown in Algorithm 5 and described as follows.
@@ 6  Randomly select a $n$-dimensional vector $\textit{{V}}^{n}$ and a number $p$;
10 Decrypt $p\times\llbracket h^{k}_{m,t}\rrbracket_{Pkv}$and $\textit{{V}}^{n}\odot\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}$ with SkvSkv;
@@ 12  Compute $\textit{Sum}^{k}_{t}$;
13  Encrypt $j^{k}_{m,t}$, $\textit{{V}}^{n}\odot\widetilde{\textit{{c}}}^{k}_{m,t}$ using ClientsPkxPkx;
@@ 19  Distribute global prototype $\llbracket\textit{{C}}^{k}_{t+1}\rrbracket_{Pkx}$ to clients;
$\llbracket\textit{{C}}^{\prime k}_{t+1}\rrbracket_{Pkv}=\frac{1}{|\mathcal{S}|}\sum_{m\in\mathcal{S}}\llbracket\textit{{c}}_{m,t}^{k}\rrbracket_{Pkv},\forall k\in\mathcal{I},$
where $\llbracket\textit{{C}}^{\prime k}_{t+1}\rrbracket_{Pkv}$ denotes the trusted prototype for class $k$ in $(t+1)$-th communication round. $\textit{{C}}^{\prime k}_{t+1}$ is considered a trusted prototype for two reasons: (i) In terms of magnitude, this is because the normalized malicious prototypes do not affect the magnitude of trusted prototype, only the direction of trusted prototype. (ii) In terms of direction, the average of all prototypes remains a plausible direction due to the small proportion of malicious clients. Subsequently, Aggregator obtains the plaintext $\|\textit{{C}}^{\prime k}_{t+1}\|$ for subsequent cosine similarity computation. To achieve the plaintext, it computes the inner product $\llbracket\textit{{C}}^{\prime k}_{t+1}\rrbracket_{Pkv}\cdot\llbracket\textit{{C}}^{\prime k}_{t+1}\rrbracket_{Pkv}$ and sends it to the Verifier for decryption, then receives the decrypted result. Then, cosine similarity, a widely used metric for measuring the angle between two vectors, is employed to evaluate the credibility of local prototype. If the direction of local prototype is similar to that of trusted prototype, it is assigned higher credibility. Formally, Aggregator computes the credibility of local prototype by the following formula:
$\begin{aligned} \llbracket\textit{sim}^{k}_{m,t}\rrbracket_{Pkv}&=\llbracket\textit{sim}(\widetilde{\textit{{c}}}^{k}_{m,t},\textit{{C}}^{\prime k}_{t+1})\rrbracket_{Pkv}=\left\llbracket\frac{\widetilde{\textit{{c}}}^{k}_{m,t}\cdot\textit{{C}}^{\prime k}_{t+1}}{\|\widetilde{\textit{{c}}}^{k}_{m,t}\|\|\textit{{C}}^{\prime k}_{t+1}\|}\right\rrbracket_{Pkv}\\ &=\frac{1}{\|\textit{{C}}^{\prime k}_{t+1}\|}\left(\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}\cdot\llbracket\textit{{C}}^{\prime k}_{t+1}\rrbracket_{Pkv}\right),\end{aligned}$
where $\llbracket\textit{sim}^{k}_{m,t}\rrbracket_{Pkv}$ denotes the credibility of local prototype. Additionally, we define a detection threshold $\chi$ and classify local prototypes with credibility less than $\chi$ as anomalous local prototypes. During global prototype aggregation, we set the aggregation weight of anomalous local prototype to 0. Formally, the aggregation weight of each local prototype is calculated as follows:
@@
where $j^{k}_{m,t}$ denotes the aggregation weight of local prototype for class $k$ at client $m$ in the $t$-th communication round. However, Aggregator cannot directly compare $\llbracket\textit{sim}^{k}_{m,t}\rrbracket_{Pkv}$ with $\chi$ within the encryption domain.
To implement comparison within encryption domain, the work in [51] offers a ciphertext comparison method that outputs the maximum value of homomorphic ciphertexts corresponding to two plaintexts in the range (0, 1) without decryption. However, since cosine similarity ranges from -1 to 1, the method cannot be applied directly. To resolve the range mismatch, Aggregator sets $\llbracket\textit{sim}^{\prime k}_{m,t}\rrbracket=\frac{1}{2}(\llbracket\textit{sim}^{k}_{m,t}\rrbracket+\llbracket 1\rrbracket)$ and $\chi^{\prime}=\frac{1}{2}(\chi+1)$ as input to the comparison method [51]. This is because the size relationship between $\llbracket\textit{sim}^{k}_{m,t}\rrbracket$ and $\chi$ remains unchanged. Then, Aggregator employs Algorithm 6 to calculate the maximum value $\llbracket h^{k}_{m,t}\rrbracket$ between $\llbracket\textit{sim}^{\prime k}_{m,t}\rrbracket$ and $\llbracket\chi^{\prime}\rrbracket$. Notably, in the Algorithm 6, $a$ and $b$ are defined as $a=\llbracket\textit{sim}^{\prime k}_{m,t}\rrbracket$ and $b=\llbracket\chi^{\prime}\rrbracket$. From these definitions, $q_{1}=\frac{a+b}{2}$ and $q_{2}=\frac{a-b}{2}$ are directly computed. The parameter $d$ in Algorithm 6 is a hyper-parameter. The larger the value of $d$, the more accurate the output. Based on previous work [51], setting $d=29$ is sufficient to ensure that the output matches the top-16 bits coincide with those of the true maximum value. In addition, the design principle process of Algorithm 6 is detailed in works [51][52].
@@
1 Initialize $a=\llbracket\textit{sim}^{\prime k}_{m,t}\rrbracket,b=\llbracket\chi^{\prime}\rrbracket$;